OpenStudy (anonymous):

Write an equation for the problem and solve. There are 17 2/3 gallons of water in the tank. The tank is 3/4 full. How many gallons of water, G, can the tank hold? My answer was G= 13 1/4. I did the inverse and got 17 2/3 gallons back that's correct.. But why is it that the tank holds less the the amount currently in it... Where did I go wrong.

OpenStudy (mertsj):

3/4x=17 2/3

OpenStudy (anonymous):

Mertsj, I am trying to figure out where did I go wrong finding G or in your case X

OpenStudy (mertsj):

\[x=\frac{53}{3}\times\frac{4}{3}=\frac{212}{9}\]

OpenStudy (mertsj):

What equation did you write?

OpenStudy (mertsj):

The thought process is; 3/4 of a full tank is 17 2/3

OpenStudy (anonymous):

\[\alpha \times 3/4= \]

OpenStudy (anonymous):

METSj Thanks!!

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

The amount the tank can hold should have been 23.55. \[(3/4) t = 17.66\] \[(1/4) t = (17.66)/4\] \[t = 17.666 *(4/3)\] \[t = 17.55\]